OpenStudy (anonymous):

Show by means of an example that lim x-> a [f(x) + g(x)] may exist evn though neither lim x-> a f(x) adn lim x -> a g(x) exists.

OpenStudy (jamesj):

Sure. Take a = 0. What's a function that doesn't have a limit there? One such function is f(x) = 1/x. Another is g(x) = x - 1/x. But f(x) + g(x) = x does have a limit at 0.

OpenStudy (zarkon):

\[f(x)=\left\{\begin{matrix}1, & x\ge 0, \\ 0 & x<0\end{matrix}\right.\] \[g(x)=\left\{\begin{matrix}0, & x\ge 0, \\ 1 & x<0\end{matrix}\right.\]

OpenStudy (anonymous):

is taht the answer? the two values of f(x) and g(x)

OpenStudy (anonymous):

also Show by means of an example that lim x-> a [f(x)g(x)] may exist evn though neither lim x-> a f(x) adn lim x -> a g(x) exists.

OpenStudy (zarkon):

I guess I should write \[f(x)=\left\{\begin{matrix}1, & x\ge a \\ 0, & x<a\end{matrix}\right.\] \[g(x)=\left\{\begin{matrix}0, & x\ge a \\ 1, & x<a\end{matrix}\right.\]

OpenStudy (zarkon):

you can use my f,g for both questions

OpenStudy (zarkon):

f+g=1 f*g=0 but neither limit exists on its own

OpenStudy (anonymous):

i see...but y is taht it doesnt exist on its own

OpenStudy (zarkon):

limit from one side is 1 and from the other side is 0...they are different so the limit does not exist

OpenStudy (anonymous):

ohhh i always forget taht if teh RHL and teh LHL do not match it does not exist

OpenStudy (anonymous):

can u help me w/ a dfft limit problem

OpenStudy (anonymous):

i dont understand it

OpenStudy (zarkon):

add a new post. I'm sure someone will help you

OpenStudy (anonymous):

ok